OpenStudy (anonymous):

A system consist of N no. of particles N>>1.Each particle can have only one of the two energies E1 or E1+ε (ε>0).If the system is in equllibrium at temp T, the avg NO.of particles with energy E1 is……

OpenStudy (michele_laino):

hint: using the \(Boltzmann\) factor, we can write the subsequent expression for the requested average number of particles: \[\Large \begin{gathered} Np = \frac{{N{e^{ - {E_1}/kT}}}}{{{e^{ - {E_1}/kT}} + {e^{ - \left( {{E_1} + \varepsilon } \right)/kT}}}} = \hfill \\ \hfill \\ = N{e^{ - {E_1}/kT}}\frac{1}{{{e^{ - {E_1}/kT}}\left( {1 + {e^{ - \varepsilon /kT}}} \right)}} = ...? \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

where \(p\) is the probability to get the system in the \(E_1\) quantum state, namely the quantum state, with energy \(E_1\)